Assume there is a file which consists of 20 blocks. Consider that the file-control block is already stored in the memory. Calculate the number of disk I/O operations are required for contiguous and linked allocation strategies, if, for one block, the following conditions hold. Also assume that the block information to be added is stored in memory.

a. The block is added at the beginning.

b. The block is added in the middle.

c. The block is added at the end.

Note: Make suitable diagrams to illustrate the complete process rather than writing the direct results or answers.
Answer #1

Answer : For evaluating the 5^532 mod 36 by using the method of repeated squaring the approach to do this is repeatedly observing the results of 5^532 mod 36 and as the result becomes continuous.

So, 5^532 mod 36 = (5^5 + 5^3 +5^2)^36 mod 0 = (3125 + 125 + 25) ^36 mod 0 = (3125^4 + 123^2 + 25^1) mod 36 = (71452)^x where x is almost reaching infinity. So as per the calculations and algorithm the answer would not be definite as per the laws of repeated squaring method.